It is a construction management problemAs a contractor you have built a 300-unit apartment complex that rents for S 1,100 per unit a month (assume 30 days/month). You were late in completing the project by 3 days and you assessed (i.e., asked to pay the owner) S 35,000. Would you call the assess s or a penalty? (3pt) Explain the basis for your answer and show your calculations. (5pt) ment liquidated damage If the contract had included a bonus/penalty clause of $50,000 per day for early completion, would you expect to gain/lose any funds had you completed the work early by 2 days? (2pt) Explain and show your calculations. Bpt)
Answer #1

All units in dollars

Per month rental = per unit rent * no of units = 1100*300= 330000

Per day rental = Per month rental / no of days in a month = 330000/30 = $ 11000

Since project is delayed by 3 days, total opportunity loss to owner = 3* $ 11000= $33000

The owner has demanded $ 35000 which is higher than the opportunity loss of $ 33000.

These can best be termed as penalty for delayed project execution (there is no liquidation or any damage involved here)

If the work had been completed 2 days earlier, the contractor would have gained $ 50000 * 2 = $ 100000